# Acylindrical hyperbolicity and the centers of Artin groups that are not free of infinity

### v2

@@ 3. Reduced clique-cube complexes and actions on them
For $i\in\{1,\ldots,k\}$, $l\in\{1,\ldots,n\}$, $\mathfrak{j}\in\{1,\ldots,\mathfrak{m}\}$, and $a\in\{1,\ldots,r\}$, we have the following:
@@
    $H_{i,l(i_{a},i_{a+1})}\cap\lambda_{l(i_{a},i_{a+1})}(i_{a},i_{a+1})H_{i,l(i_{a},i_{a+1})}=\emptyset$;
@@
    $H_{i,l(w_{\mathfrak{j}})}\cap\lambda_{l(w_{\mathfrak{j}})}(\mathfrak{j})H_{i,l(w_{\mathfrak{j}})}=\emptyset$;
@@
    $[A_{V_{0}},A_{U_{l(i_{a},i_{a+1})}}]\cap\lambda_{l(i_{a},i_{a+1})}(i_{a},i_{a+1})[A_{V_{0}},A_{U_{l(i_{a},i_{a+1})}}]=\{A_{U_{l(i_{a},i_{a+1})}}\}$.
@@
    $[A_{V_{0}},A_{U_{l(w_{\mathfrak{j}})}}]\cap\lambda_{l(w_{\mathfrak{j}})}(\mathfrak{j})[A_{V_{0}},A_{U_{l(w_{\mathfrak{j}})}}]=\{A_{U_{l(w_{\mathfrak{j}})}}\}$.
@@
(1) $H_{i,l}$ and $H_{i,l+1}$ are of $v_{i,l}$-type and $v_{i,l+1}$-type, respectively. Note that $v_{i,l}\neq v_{i,l+1}$ implies $H_{i,l}\neq H_{i,l+1}$. Because $(v_{i,l},v_{i,l+1})\notin E(\Gamma)$, we have $\{v_{i,l},v_{i,l+1}\}\notin\Delta_{\Gamma}$, and thus $H_{i,l}\cap H_{i,l+1}=\emptyset$.
(2) Because $[A_{U_{l}\setminus\{v_{i,l}\}},A_{U_{l}}]\subset N(H_{i,l})$, we have $[\lambda_{l}A_{U_{l}\setminus\{v_{i,l}\}},\lambda_{l}A_{U_{l}}]\subset\lambda_{l}N(H_{i,l})$. Note that $\lambda_{l}A_{U_{l}}=A_{U_{l}}$ and $\lambda_{l}A_{U_{l}\setminus\{v_{i,l}\}}=v_{1,l}\cdots v_{i,l}A_{U_{l}\setminus\{v_{i,l}\}}$. Now, assume that $H_{i,l}\cap\lambda_{l}H_{i,l}\neq\emptyset$. Because $H_{i,l}$ and $\lambda_{l}H_{i,l}$ are of $v_{i,l}$-type, we see that $H_{i,l}=\lambda_{l}H_{i,l}$ (see Remark 2.13). Then, we have
@@
Hence, we obtain $A_{U_{l}\setminus\{v_{i,l}\}}=v_{1,l}\cdots v_{i,l}A_{U_{l}\setminus\{v_{i,l}\}}$. This means that $v_{i,l}\in A_{U_{l}\setminus\{v_{i,l}\}}$. However, from Lemma 2.10 (2), $\{v_{i,l}\}\cap(U_{l}\setminus\{v_{i,l}\})=\emptyset$ implies that $A_{\{v_{i,l}\}}\cap A_{U_{l}\setminus\{v_{i,l}\}}=A_{\emptyset}=\{1\}$. This contradicts $v_{i,l}\neq 1$ in $A_{\Gamma}$.
@@
$H_{i,l(i_{a},i_{a+1})}\cap\lambda_{l(i_{a},i_{a+1})}(i_{a},i_{a+1})H_{i,l(i_{a},i_{a+1})}=\emptyset.$
@@
$H_{i,l(i_{a},i_{a+1})}\cap\lambda_{l(i_{a},i_{a+1})}(i_{a},i_{a+1})H_{i,l(i_{a},i_{a+1})}\neq\emptyset.$
Note that $\lambda_{l(i_{a},i_{a+1})}(i_{a},i_{a+1})A_{U_{l(i_{a},i_{a+1})}}=A_{U_{l(i_{a},i_{a+1})}}$ and
$\lambda_{l(i_{a},i_{a+1})}(i_{a},i_{a+1})A_{U_{l(i_{a},i_{a+1})}\setminus\{v_{i,l(i_{a},i_{a+1})}\}}=\tau_{i_{a},i_{a+1}}A_{U_{l(i_{a},i_{a+1})}\setminus\{v_{i,l(i_{a},i_{a+1})}\}}.$
Because $H_{i,l(i_{a},i_{a+1})}$ and $\lambda_{l(i_{a},i_{a+1})}H_{i,l(i_{a},i_{a+1})}$ are of $v_{i,l(i_{a},i_{a+1})}$-type, we see that $H_{i,l(i_{a},i_{a+1})}=\lambda_{l(i_{a},i_{a+1})}H_{i,\lambda_{l(i_{a},i_{a+1})}}$. Then, we have
$\begin{split}[\lambda_{l(i_{a},i_{a+1})}A_{U_{l(i_{a},i_{a+1})}\setminus\{v_{i,l(i_{a},i_{a+1})}\}},\lambda_{l(i_{a},i_{a+1})}A_{U_{l(i_{a},i_{a+1})}}]\\ =[\tau_{i_{a},i_{a+1}}A_{U_{l(i_{a},i_{a+1})}\setminus\{v_{i,{l(i_{a},i_{a+1})}}\}},A_{U_{l(i_{a},i_{a+1})}}]\subset N(H_{i,{l(i_{a},i_{a+1})}}).\end{split}$
@@
$\tau_{i_{a},i_{a+1}}A_{U_{l(i_{a},i_{a+1})}\setminus\{v_{i,l(i_{a},i_{a+1})}\}}=A_{U_{l(i_{a},i_{a+1})}\setminus\{v_{i,l(i_{a},i_{a+1})}\}}.$
@@
The following is a variant for $C_{\Gamma,\Gamma_{0}}$ of [29, Lemma 6.3] for $C_{\Gamma}$, and its proof is same.
@@
$H_{i_{a},l(i_{a},i_{a+1})}\cap\lambda_{l(i_{a},i_{a+1})}(i_{a},i_{a+1})H_{i_{a+1},l(i_{a},i_{a+1})}=\emptyset.$
@@
$\ldots,J_{i,-1},J_{i,0},J_{i,1},\ldots,J_{i,2(\mathfrak{m}+r)n},J_{i,2(\mathfrak{m}+r)n+1},\ldots,$
a sequence of $k$-dimensional cubes
$\ldots,K_{-1},K_{0},K_{1},\cdots,K_{2(\mathfrak{m}+r)n},K_{2(\mathfrak{m}+r)n+1},\ldots,$
and a sequence of vertices of $C_{\Gamma,\Gamma_{0}}$
$\ldots,w_{-1},w_{0},w_{1},\ldots,w_{2(\mathfrak{m}+r)n},w_{2(\mathfrak{m}+r)n+1},\ldots$
as follows. First, for $a\in\{1,\ldots,\mathfrak{m}+r\}$ and $l\in\{1,\dots,n\}$, we define
$J_{i,2((a-1)n+l)-1}$ $:=\gamma((a-1)n+l-1)H_{i,l},$
$J_{i,2((a-1)n+l)}$ $:=\gamma((a-1)n+l)H_{i,l},$
$K_{2((a-1)n+l)-1}$ $:=\gamma((a-1)n+l-1)[A_{V_{0}},A_{U_{l}}],$
$K_{2((a-1)n+l)}$ $:=\gamma((a-1)n+l)[A_{V_{0}},A_{U_{l}}],$
$w_{2((a-1)n+l)-1}$ $:=\gamma((a-1)n+l-1)A_{U_{l}},$
$w_{2((a-1)n+l)}$ $:=\gamma((a-1)n+l)A_{V_{0}},$
where we note that both $J_{i,2((a-1)n+l)-1}$ and $J_{i,2((a-1)n+l)}$ are of the same $v_{i,l}$-type.
Second, for any $b\in\{1,\dots,2(\mathfrak{m}+r)n\}$ and $c\in\mathbb{Z}$, we set
$J_{i,2(\mathfrak{m}+r)nc+b}:=\gamma^{c}J_{i,b},K_{2(\mathfrak{m}+r)nc+b}:=\gamma^{c}K_{b},w_{2(\mathfrak{m}+r)nc+b}:=\gamma^{c}w_{b}.$
Then, we have
$K_{2(\mathfrak{m}+r)nc+b}=[w_{2(\mathfrak{m}+r)nc+b-1},w_{2(\mathfrak{m}+r)nc+b}].$
Additionally, we have the two connected components $J_{i,2(\mathfrak{m}+r)nc+b}^{-}$ and $J_{i,2(\mathfrak{m}+r)nc+b}^{+}$ such that
$C_{\Gamma}\!\setminus\!\!\setminus J_{i,2(\mathfrak{m}+r)nc+b}=J_{i,2(\mathfrak{m}+r)nc+b}^{-}\sqcup J_{i,2(\mathfrak{m}+r)nc+b}^{+},$
$w_{2(\mathfrak{m}+r)nc+b-1}\in J_{i,2(\mathfrak{m}+r)nc+b}^{-},w_{2(\mathfrak{m}+r)nc+b}\in J_{i,2(\mathfrak{m}+r)nc+b}^{+}.$
@@
$\cdots\subsetneq J_{i,-1}^{-}\subsetneq J_{i,0}^{-}\subsetneq J_{i,1}^{-}\subsetneq\cdots\subsetneq J_{i,2(\mathfrak{m}+r)n}^{-}\subsetneq J_{i,2(\mathfrak{m}+r)n+1}^{-}\subsetneq\cdots,$
$\cdots\supsetneq J_{i,-1}^{+}\supsetneq J_{i,0}^{+}\supsetneq J_{i,1}^{+}\supsetneq\cdots\supsetneq J_{i,2(\mathfrak{m}+r)n}^{+}\supsetneq J_{i,2(\mathfrak{m}+r)n+1}^{+}\supsetneq\cdots.$
@@
$J_{i,0}^{-}\not\ni w_{0}=A_{V_{0}}\in J_{i,1}^{-},$
$J_{i,2(\mathfrak{m}+r)n}^{+}\ni w_{2(\mathfrak{m}+r)n}=\gamma A_{V_{0}}\not\in J_{i,2(\mathfrak{m}+r)n+1}^{+}.$
We have the following, which is a variant for $C_{\Gamma,\Gamma_{0}}$ of [29, Lemma 6.4] for $C_{\Gamma}$.
@@
The set $\{J_{i,d}\}_{i\in\{1,\ldots,k\},d\in\{1,\ldots,2(\mathfrak{m}+r)n\}}$ is the set of all hyperplanes separating $w_{0}=A_{V_{0}}$ and $w_{2(\mathfrak{m}+r)n}=\gamma A_{V_{0}}$.
For the proof, we consider a path from $w_{0}=A_{V_{0}}$ to $w_{2(\mathfrak{m}+r)n}=\gamma A_{V_{0}}$ that diagonally penetrates each of the cubes $K_{1},\ldots,K_{2(\mathfrak{m}+r)n}$ in order. The set of all hyperplanes intersecting the path is $\{J_{i,d}\}_{i\in\{1,\ldots,k\},d\in\{1,\ldots,2(\mathfrak{m}+r)n\}}$.
We now state the final lemma required for the proof of Theorem 1.2, where we recall Definition 2.7 and Remark 2.8. The following is a variant for $C_{\Gamma,\Gamma_{0}}$ of [29, Lemma 6.5] for $C_{\Gamma}$.
@@
    $J_{i,1},\ldots,J_{i,2(\mathfrak{m}+r)n}$
    is a sequence of separating hyperplanes from $J_{i,0}$ to $J_{i,2(\mathfrak{m}+r)n+1}$, where $J_{i,0}$ and $J_{i,2(\mathfrak{m}+r)n+1}$ are of $v_{i,n}$-type and $v_{i,1}$-type, respectively (in particular, from $w_{0}=A_{V_{0}}$ to $w_{2(\mathfrak{m}+r)n}=\gamma A_{V_{0}}$).
@@
    $J_{i_{1},1},J_{i_{1},2},\ldots,J_{i_{1},2\mathfrak{m}n},$
    $J_{i_{1},2\mathfrak{m}n+1},J_{i_{1},2\mathfrak{m}n+2},\ldots,J_{i_{1},2\mathfrak{m}n+2l(i_{1},i_{2})-1},$
    $J_{i_{2},2\mathfrak{m}n+2l(i_{1},i_{2})},J_{i_{2},2\mathfrak{m}n+2l(i_{1},i_{2})+1},\ldots,J_{i_{2},2\mathfrak{m}n+2n},$
    $J_{i_{2},2\mathfrak{m}n+2n+1},J_{i_{2},2\mathfrak{m}n+2(n+1)},\ldots,J_{i_{2},2\mathfrak{m}n+2(n+l(i_{2},i_{3}))-1},$
    $\cdots$
    $J_{i_{r},2\mathfrak{m}n+2((r-2)n+l(i_{r-1},i_{r}))},J_{i_{r},2\mathfrak{m}n+2((r-2)n+l(i_{r-1},i_{r}))+1},\ldots,J_{i_{r},2\mathfrak{m}n+2(r-1)n},$
    $J_{i_{r},2\mathfrak{m}n+2(r-1)n+1},J_{i_{r},2\mathfrak{m}n+2((r-1)n+1)},\ldots,J_{i_{r},2\mathfrak{m}n+2((r-1)n+l(i_{r},i_{r+1}))-1},$
    $J_{i_{r+1},2\mathfrak{m}n+2((r-1)n+l(i_{r},i_{r+1}))},J_{i_{r+1},2\mathfrak{m}n+2((r-1)n+l(i_{r},i_{r+1}))+1},\ldots,J_{i_{r+1},2(\mathfrak{m}+r)n}$
    is a sequence of separating hyperplanes from $J_{1,0}$ to $J_{1,2(\mathfrak{m}+r)n+1}$, where $J_{1,0}$ and $J_{1,2(\mathfrak{m}+r)n+1}$ are of $v_{1,n}$-type and $v_{1,1}$-type, respectively (in particular, from $w_{0}=A_{V_{0}}$ to $w_{2(\mathfrak{m}+r)n}=\gamma A_{V_{0}}$). Here note that $i_{1}=1$. (See Figure 5.)
    Moreover, the sequence contains a hyperplane of $v_{i}$-type for any $i\in\{1,\ldots,k\}$ and $v_{i}\in V(\Gamma_{i})$.
Figure 5. Part of a sequence of hyperplanes for the case where $\Gamma=\Gamma_{0}\ast\Gamma_{1}\ast\Gamma_{2}$.

### v3

@@ 3. Reduced clique-cube complexes and actions on them
For $i\in\{1,\ldots,k\}$, $l\in\{1,\ldots,n\}$, $\mathfrak{j}\in\{1,\ldots,\mathfrak{m}\}$ and $a\in\{1,\ldots,r\}$, we have the following:
@@
    $H_{i,l_{\flat}(i_{a},i_{a+1})}\cap\lambda_{\flat}(a)H_{i,l_{\flat}(i_{a},i_{a+1})}=\emptyset$;
@@
    $H_{i,l_{\sharp}(w_{\mathfrak{j}})}\cap\lambda_{\sharp}(\mathfrak{j})H_{i,l_{\sharp}(w_{\mathfrak{j}})}=\emptyset$;
@@
    $[A_{V_{0}},A_{U_{l_{\flat}(i_{a},i_{a+1})}}]\cap\lambda_{\flat}(a)[A_{V_{0}},A_{U_{l_{\flat}(i_{a},i_{a+1})}}]=\{A_{U_{l_{\flat}(i_{a},i_{a+1})}}\}$.
@@
    $[A_{V_{0}},A_{U_{l_{\sharp}(w_{\mathfrak{j}})}}]\cap\lambda_{\sharp}(\mathfrak{j})[A_{V_{0}},A_{U_{l_{\sharp}(w_{\mathfrak{j}})}}]=\{A_{U_{l_{\sharp}(w_{\mathfrak{j}})}}\}$.
@@
(1) $H_{i,l}$ and $H_{i,l+1}$ are of $v_{i,l}$-type and $v_{i,l+1}$-type, respectively. Note that $v_{i,l}\neq v_{i,l+1}$ implies $H_{i,l}\neq H_{i,l+1}$. Because $(v_{i,l},v_{i,l+1})\notin E$, we have $H_{i,l}\cap H_{i,l+1}=\emptyset$ (see Remark 2.14).
(2) Because $[A_{U_{l}\setminus\{v_{i,l}\}},A_{U_{l}}]\subset N(H_{i,l})$, we have $[\lambda_{l}A_{U_{l}\setminus\{v_{i,l}\}},\lambda_{l}A_{U_{l}}]\subset\lambda_{l}N(H_{i,l})$. Note that $\lambda_{l}A_{U_{l}}=A_{U_{l}}$ and $\lambda_{l}A_{U_{l}\setminus\{v_{i,l}\}}=v_{1,l}\cdots v_{i,l}A_{U_{l}\setminus\{v_{i,l}\}}$. Assume that $H_{i,l}\cap\lambda_{l}H_{i,l}\neq\emptyset$. Because $H_{i,l}$ and $\lambda_{l}H_{i,l}$ are of $v_{i,l}$-type, we see that $H_{i,l}=\lambda_{l}H_{i,l}$ (see Remark 2.14). Then, we have
@@
Hence, we obtain $A_{U_{l}\setminus\{v_{i,l}\}}=v_{1,l}\cdots v_{i,l}A_{U_{l}\setminus\{v_{i,l}\}}$. Hence, $A_{U_{l}\setminus\{v_{i,l}\}}=v_{i,l}A_{U_{l}\setminus\{v_{i,l}\}}$. This means that $v_{i,l}\in A_{U_{l}\setminus\{v_{i,l}\}}$. However, from Lemma 2.10 (2), $\{v_{i,l}\}\cap(U_{l}\setminus\{v_{i,l}\})=\emptyset$ implies that $A_{\{v_{i,l}\}}\cap A_{U_{l}\setminus\{v_{i,l}\}}=A_{\emptyset}=\{e\}$. This contradicts $v_{i,l}\neq e$ in $A_{\Gamma}$.
@@
$H_{i,l_{\flat}(i_{a},i_{a+1})}\cap\lambda_{\flat}(a)H_{i,l_{\flat}(i_{a},i_{a+1})}=\emptyset.$
@@
$H_{i,l_{\flat}(i_{a},i_{a+1})}\cap\lambda_{\flat}(a)H_{i,l_{\flat}(i_{a},i_{a+1})}\neq\emptyset.$
Note that $\lambda_{\flat}(a)A_{U_{l_{\flat}(i_{a},i_{a+1})}}=A_{U_{l_{\flat}(i_{a},i_{a+1})}}$ and
$\lambda_{\flat}(a)A_{U_{l_{\flat}(i_{a},i_{a+1})}\setminus\{v_{i,l_{\flat}(i_{a},i_{a+1})}\}}=\tau_{i_{a},i_{a+1}}A_{U_{l_{\flat}(i_{a},i_{a+1})}\setminus\{v_{i,l_{\flat}(i_{a},i_{a+1})}\}}.$
Because both $H_{i,l_{\flat}(i_{a},i_{a+1})}$ and $\lambda_{\flat}(a)H_{i,l_{\flat}(i_{a},i_{a+1})}$ are of $v_{i,l_{\flat}(i_{a},i_{a+1})}$-type, we see that $H_{i,l_{\flat}(i_{a},i_{a+1})}=\lambda_{\flat}(a)H_{i,l_{\flat}(i_{a},i_{a+1})}$. Then, we have
$\begin{split}[\lambda_{\flat}(a)A_{U_{l_{\flat}(i_{a},i_{a+1})}\setminus\{v_{i,l_{\flat}(i_{a},i_{a+1})}\}},\lambda_{\flat}(a)A_{U_{l_{\flat}(i_{a},i_{a+1})}}]\\ =[\tau_{i_{a},i_{a+1}}A_{U_{l_{\flat}(i_{a},i_{a+1})}\setminus\{v_{i,{l_{\flat}(i_{a},i_{a+1})}}\}},A_{U_{l_{\flat}(i_{a},i_{a+1})}}]\subset N(H_{i,{l_{\flat}(i_{a},i_{a+1})}}).\end{split}$
@@
$\tau_{i_{a},i_{a+1}}A_{U_{l_{\flat}(i_{a},i_{a+1})}\setminus\{v_{i,l_{\flat}(i_{a},i_{a+1})}\}}=A_{U_{l_{\flat}(i_{a},i_{a+1})}\setminus\{v_{i,l_{\flat}(i_{a},i_{a+1})}\}}.$
@@
The following is a variant for $C_{\Gamma,\Gamma_{0}}$ of [29, Lemma 6.3] for $C_{\Gamma}$, and its proof is the same. In fact, Lemma 3.9 plays the same role that [29, Lemma 5.2] did in the proof of [29, Lemma 6.3].
@@
$H_{i_{a},l_{\flat}(i_{a},i_{a+1})}\cap\lambda_{\flat}(a)H_{i_{a+1},l_{\flat}(i_{a},i_{a+1})}=\emptyset.$
Figure 8. All hyperplanes in Lemma 4.6 for the case where $\Gamma=\Gamma_{0}\ast\Gamma_{1}\ast\Gamma_{2}$. Hyperplanes $J_{1,d}$ and $J_{2,d}$ intersecting $K_{d}$ are represented as the solid and the dotted lines, respectively.
@@
$\ldots,J_{i,-1},J_{i,0},J_{i,1},\ldots,J_{i,2(1+\mathfrak{m}+r)n},J_{i,2(1+\mathfrak{m}+r)n+1},\ldots,$
and a sequence of $k$-dimensional cubes
$\ldots,K_{-1},K_{0},K_{1},\cdots,K_{2(1+\mathfrak{m}+r)n},K_{2(1+\mathfrak{m}+r)n+1},\ldots$
as follows. First, for $a\in\{1,\ldots,1+\mathfrak{m}+r\}$ and $l\in\{1,\dots,n\}$, we define
$J_{i,2((a-1)n+l)-1}$ $:=\gamma_{\natural\sharp\flat}[(a-1)n+l-1]H_{i,l},$
$J_{i,2((a-1)n+l)}$ $:=\gamma_{\natural\sharp\flat}[(a-1)n+l]H_{i,l},$
$K_{2((a-1)n+l)-1}$ $:=\gamma_{\natural\sharp\flat}[(a-1)n+l-1][A_{V_{0}},A_{U_{l}}]$
$=[\gamma_{\natural\sharp\flat}[(a-1)n+l-1]A_{V_{0}},\gamma_{\natural\sharp\flat}[(a-1)n+l-1]A_{U_{l}}],$
$K_{2((a-1)n+l)}$ $:=\gamma_{\natural\sharp\flat}[(a-1)n+l][A_{V_{0}},A_{U_{l}}]$
$=[\gamma_{\natural\sharp\flat}[(a-1)n+l]A_{V_{0}},\gamma_{\natural\sharp\flat}[(a-1)n+l]A_{U_{l}}]$
$=[\gamma_{\natural\sharp\flat}[(a-1)n+l]A_{V_{0}},\gamma_{\natural\sharp\flat}[(a-1)n+l-1]A_{U_{l}}].$
Here, we note that both $J_{i,2((a-1)n+l)-1}$ and $J_{i,2((a-1)n+l)}$ are of the same $v_{i,l}$-type, and that we have $\gamma_{\natural\sharp\flat}[(a-1)n+l-1]A_{U_{l}}=\gamma_{\natural\sharp\flat}[(a-1)n+l]A_{U_{l}}$. See Figure 8.
Second, for any $b\in\{1,\dots,2(1+\mathfrak{m}+r)n\}$ and $c\in\mathbb{Z}$, we set
$J_{i,2(1+\mathfrak{m}+r)nc+b}:={(\gamma_{\natural\sharp\flat})}^{c}J_{i,b},$
$K_{2(1+\mathfrak{m}+r)nc+b}:={(\gamma_{\natural\sharp\flat})}^{c}K_{b}.$
Additionally, we have the two connected components $J_{i,2(1+\mathfrak{m}+r)nc+b}^{-}$ and $J_{i,2(1+\mathfrak{m}+r)nc+b}^{+}$ such that
$C_{\Gamma}\!\setminus\!\!\setminus J_{i,2(1+\mathfrak{m}+r)nc+b}=J_{i,2(1+\mathfrak{m}+r)nc+b}^{-}\sqcup J_{i,2(1+\mathfrak{m}+r)nc+b}^{+},$
and the “left endpoint” of $K_{2(1+\mathfrak{m}+r)nc+b}$ is in $J_{i,2(1+\mathfrak{m}+r)nc+b}^{-}$ while the “right endpoint” of $K_{2(1+\mathfrak{m}+r)nc+b}$ is in $J_{i,2(1+\mathfrak{m}+r)nc+b}^{+}$. Precisely, we have
${(\gamma_{\natural\sharp\flat})}^{c}\left(\gamma_{\natural\sharp\flat}\left[\frac{b-1}{2}\right]\right)A_{V_{0}}\in J_{i,2(1+\mathfrak{m}+r)nc+b}^{-},$
${(\gamma_{\natural\sharp\flat})}^{c}\left(\gamma_{\natural\sharp\flat}\left[\frac{b-1}{2}\right]\right)A_{U_{l}}\in J_{i,2(1+\mathfrak{m}+r)nc+b}^{+}$
if $b=2((a-1)n+l)-1$ with some $a\in\{1,\ldots,1+\mathfrak{m}+r\}$ and $l\in\{1,\dots,n\}$, and we have
${(\gamma_{\natural\sharp\flat})}^{c}\left(\gamma_{\natural\sharp\flat}\left[\frac{b}{2}\right]\right)A_{U_{l}}\in J_{i,2(1+\mathfrak{m}+r)nc+b}^{-},$
${(\gamma_{\natural\sharp\flat})}^{c}\left(\gamma_{\natural\sharp\flat}\left[\frac{b}{2}\right]\right)A_{V_{0}}\in J_{i,2(1+\mathfrak{m}+r)nc+b}^{+}$
if $b=2((a-1)n+l)$ with some $a\in\{1,\ldots,1+\mathfrak{m}+r\}$ and $l\in\{1,\dots,n\}$. Note that in the former case,
$K_{2(1+\mathfrak{m}+r)nc+b}$ $={(\gamma_{\natural\sharp\flat})}^{c}\left(\gamma_{\natural\sharp\flat}\left[\frac{b-1}{2}\right]\right)[A_{V_{0}},A_{U_{l}}]$
$=\left[{(\gamma_{\natural\sharp\flat})}^{c}\left(\gamma_{\natural\sharp\flat}\left[\frac{b-1}{2}\right]\right)A_{V_{0}},{(\gamma_{\natural\sharp\flat})}^{c}\left(\gamma_{\natural\sharp\flat}\left[\frac{b-1}{2}\right]\right)A_{U_{l}}\right]$
and in the latter case,
$K_{2(1+\mathfrak{m}+r)nc+b}=\left[{(\gamma_{\natural\sharp\flat})}^{c}\left(\gamma_{\natural\sharp\flat}\left[\frac{b-2}{2}\right]\right)A_{U_{l}},{(\gamma_{\natural\sharp\flat})}^{c}\left(\gamma_{\natural\sharp\flat}\left[\frac{b}{2}\right]\right)A_{V_{0}}\right].$
@@
$\cdots\subsetneq J_{i,-1}^{-}\subsetneq J_{i,0}^{-}\subsetneq J_{i,1}^{-}\subsetneq\cdots\subsetneq J_{i,2(1+\mathfrak{m}+r)n}^{-}\subsetneq J_{i,2(1+\mathfrak{m}+r)n+1}^{-}\subsetneq\cdots,$
$\cdots\supsetneq J_{i,-1}^{+}\supsetneq J_{i,0}^{+}\supsetneq J_{i,1}^{+}\supsetneq\cdots\supsetneq J_{i,2(1+\mathfrak{m}+r)n}^{+}\supsetneq J_{i,2(1+\mathfrak{m}+r)n+1}^{+}\supsetneq\cdots.$
@@
$J_{i,0}^{-}\not\ni A_{V_{0}}\in J_{i,1}^{-},$
$J_{i,2(1+\mathfrak{m}+r)n}^{+}\ni\gamma_{\natural\sharp\flat}A_{V_{0}}\not\in J_{i,2(1+\mathfrak{m}+r)n+1}^{+}.$
@@
The set $\{J_{i,d}\}_{i\in\{1,\ldots,k\},d\in\{1,\ldots,2(1+\mathfrak{m}+r)n\}}$ is the set of all hyperplanes separating $A_{V_{0}}$ and $\gamma_{\natural\sharp\flat}A_{V_{0}}$.
See Figure 8.
To prove the lemma, we consider a path $\ell$ from $A_{V_{0}}$ to $\gamma_{\natural\sharp\flat}A_{V_{0}}$ that diagonally penetrates each of the cubes $K_{1},\ldots,K_{2(1+\mathfrak{m}+r)n}$ in order. The set of all hyperplanes intersecting the path $\ell$ is $\{J_{i,d}\}_{i\in\{1,\ldots,k\},d\in\{1,\ldots,2(1+\mathfrak{m}+r)n\}}$.
We now state the final lemma required to complete the proof of Theorem 4.1. Recall Definition 2.7 and Remark 2.8 for sequences of separating hyperplanes. The following is a variant for $C_{\Gamma,\Gamma_{0}}$ of [29, Lemma 6.5] for $C_{\Gamma}$.
@@
    $J_{i,1},\ldots,J_{i,2(1+\mathfrak{m}+r)n}$
    is a sequence of separating hyperplanes from $J_{i,0}$ to $J_{i,2(1+\mathfrak{m}+r)n+1}$, where $J_{i,0}$ and $J_{i,2(1+\mathfrak{m}+r)n+1}$ are of $v_{i,n}$-type and $v_{i,1}$-type, respectively (in particular, from $A_{V_{0}}$ to $\gamma_{\natural\sharp\flat}A_{V_{0}}$). See Figure 8.
@@
      $J_{1,1},J_{1,2},\ldots,J_{1,2n}$
      is a sequence of separating hyperplanes from
      $\gamma_{\natural\sharp\flat}[0]A_{V_{0}}=A_{V_{0}}\text{ to }\gamma_{\natural\sharp\flat}[n]A_{V_{0}}=\gamma_{\natural}A_{V_{0}}.$
    2. (2-ii)
      The sequence of hyperplanes
      $J_{1,2n+1},J_{1,2n+2},\ldots,J_{1,2(1+\mathfrak{m})n}$
      is a sequence of separating hyperplanes from
      $\gamma_{\natural\sharp\flat}[n]A_{V_{0}}=\gamma_{\natural}A_{V_{0}}\text{ to }\gamma_{\natural\sharp\flat}[(1+\mathfrak{m})n]A_{V_{0}}=\gamma_{\natural}\gamma_{\sharp}A_{V_{0}}.$
    3. (2-iii-$a$)
      For any $a\in\{1,\ldots,r\}$, the sequence of hyperplanes
      $J_{i_{a},2(1+\mathfrak{m})n+2(a-1)n+1},J_{i_{a},2(1+\mathfrak{m})n+2((a-1)n+1)},\ldots,$
      $J_{i_{a},2(1+\mathfrak{m})n+2((a-1)n+l_{\flat}(i_{a},i_{a+1}))-1},J_{i_{a+1},2(1+\mathfrak{m})n+2((a-1)n+l_{\flat}(i_{a},i_{a+1}))},$
      $J_{i_{a+1},2(1+\mathfrak{m})n+2((a-1)n+l_{\flat}(i_{a},i_{a+1}))+1},\ldots,J_{i_{a+1},2(1+\mathfrak{m}+a)n}$
      is a sequence of separating hyperplanes from
      $\gamma_{\natural\sharp\flat}[(1+\mathfrak{m}+a-1)n]A_{V_{0}}\text{ to }\gamma_{\natural\sharp\flat}[(1+\mathfrak{m}+a)n]A_{V_{0}}.$
      See Figure 9.
  3. (3)
    The sequence of hyperplanes composed of the sequences in (2-i), (2-ii), (2-iii-$1$), (2-iii-$2$), $\ldots$, (2-iii-$r$), in order, is a sequence of separating hyperplanes from $J_{1,0}$ to $J_{1,2(1+\mathfrak{m}+r)n+1}$, where $J_{1,0}$ and $J_{1,2(1+\mathfrak{m}+r)n+1}$ are of $v_{1,n}$-type and $v_{1,1}$-type, respectively (in particular, from $A_{V_{0}}$ to $\gamma_{\natural\sharp\flat}A_{V_{0}}$). Here, note that $i_{1}=1$. Moreover, the sequence contains a hyperplane of $v_{i}$-type for any $i\in\{1,\ldots,k\}$ and any $v_{i}\in V_{i}$.
Figure 9. The sequence of hyperplanes (2-iii-1) in Lemma 4.7 for the case where $\Gamma=\Gamma_{0}\ast\Gamma_{1}\ast\Gamma_{2}$. The two hyperplanes $J_{i_{1},2(1+\mathfrak{m})n+2l_{\flat}(i_{1},i_{2})-1}$ and $J_{i_{2},2(1+\mathfrak{m})n+2l_{\flat}(i_{1},i_{2})}$ do not intersect each other.